OpenStudy (anonymous):

Let S be the solid bounded by the surface z = xy, the cylinders y = x^2 and y^2 = x and the plane z = 0. Find the volume.

terenzreignz (terenzreignz):

This is nauseating :D I haven't done this in a while, this involves double integrals, right?

OpenStudy (anonymous):

yes :)

terenzreignz (terenzreignz):

Well, then, it looks like you're integrating a function z = f(x,y) = xy over a region. This SHOULD be easy, the tricky part is getting the region...

OpenStudy (anonymous):

can you help me solving this?

terenzreignz (terenzreignz):

Let's put that in maths language, shall we? \[\huge \int\limits \int\limits xy \ dA\]Now all that remains is to find the region....

terenzreignz (terenzreignz):

Now, the region is the cylinder formed by the bounded region of y = x^2 and x = y^2

terenzreignz (terenzreignz):

And the plane z = 0, of course, but that's a mere formality :) Let's take a look at the graphs of y = x^2 and x = y^2 WARNING I SUCK AT DRAWING

terenzreignz (terenzreignz):

 Say this is your xy plane

terenzreignz (terenzreignz):

 This is your y = x^2, a parabola :)